Archie is fed up with waiting in line at his local post office and decides to take action. Over the course of the next few months, he records the waiting times for each of a random selection of 18 post office visits made by him and other customers. These waiting times (in minutes) are as follows:
27,26,13,16,8,23,22,15,16,21,11,25,30,18,12,5,10,24
construct a box-and-whisker plot for the data.



Answer by Mathtut(3670)

points are made under the lowest, highest,1st quartile, median, 3rd quartile
:
then a box is created around the 1st and 3rd quartiles with a line made throught the median then whiskers are drawn to the lowest and highest data.
:
lowest-5
1st quartile is 4th and 5th terms divided by 2 : 11+12/2=11.5
median is the 9th and 10 terms added and divided by 2: 16+17/2=16.5
3rd quartile is 14 and 15 terms added and divided by 2: 24+25/2=24.5
highest-30